OpenStudy (anonymous):

Carrie made a square tablecloth with an area of 169 square inches. What was the length of each side of the tablecloth? 9 inches 13 inches 17 inches 23 inches

OpenStudy (anonymous):

@amistre64 @acxbox22

OpenStudy (anonymous):

\[\sqrt{169}=13 \]

OpenStudy (anonymous):

so 13?

OpenStudy (anonymous):

@robtobey

jimthompson5910 (jim_thompson5910):

yep 13

OpenStudy (anonymous):

what about this?

jimthompson5910 (jim_thompson5910):

your thoughts on it?

OpenStudy (anonymous):

no A or B...

jimthompson5910 (jim_thompson5910):

agreed, what else?

OpenStudy (anonymous):

D is the answer!

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

oh, C!!!

jimthompson5910 (jim_thompson5910):

20 is the perimeter

jimthompson5910 (jim_thompson5910):

4 times side length = 4*5 = 20

OpenStudy (anonymous):

ok....C?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

What about this?

OpenStudy (anonymous):

what about this?

OpenStudy (anonymous):

If s is the side length of the square, then the area is: s^2 = 169 Take the square root of each side to calculate the length of s.

OpenStudy (anonymous):

thats a different question?

OpenStudy (anonymous):

No. I did not elaborate on the approach to the solution on my first posting.

OpenStudy (anonymous):

oh, its not 13?

OpenStudy (acxbox22):

it is 13